### v1

# Addressing Source Scale Bias via Image Warping for Domain Adaptation

Abstract
In visual recognition, scale bias is a key challenge due to the imbalance of object and image size distribution inherent in real scene datasets. Conventional solutions involve injecting scale invariance priors, oversampling the dataset at different scales during training, or adjusting scale at inference. While these strategies mitigate scale bias to some extent, their ability to adapt across diverse datasets is limited. Besides, they increase computational load during training and latency during inference. In this work, we use adaptive attentional processing – oversampling salient object regions by warping images in-place during training. Discovering that shifting the source scale distribution improves backbone features, we developed a instance-level warping guidance aimed at object region sampling to mitigate source scale bias in domain adaptation. Our approach improves adaptation across geographies, lighting and weather conditions, is agnostic to the task, domain adaptation algorithm, saliency guidance, and underlying model architecture. Highlights include +6.1 mAP50 for BDD100K Clear DENSE Foggy, +3.7 mAP50 for BDD100K Day Night, +3.0 mAP50 for BDD100K Clear Rainy, and +6.3 mIoU for Cityscapes ACDC. Our approach adds minimal memory during training and has no additional latency at inference time. Please see Appendix for more results and analysis.
![[Uncaptioned image]](x1.png)
1 Introduction
Visual detection and segmentation of things (i.e. object instances) and stuff, across different scales is a fundamental challenge in computer vision [49, 35, 19, 45]. Natural scene datasets used for training different tasks often exhibit bias of imaged objects towards certain scales (or sizes). For instance, in autonomous driving situations, camera perspective and field of view warrant that many more distant objects that appear small in the image are captured compared to nearby objects that appear large, even if the real object 3D scale distribution is near-uniform. This results in an inherent imbalance or bias in the distribution of sizes of object projections in the image, leading to sub-optimal recognition models.
Common approaches to address scale bias are: (a) selecting an image subset to create a uniform scale distribution but this is suboptimal as only a fraction of the available data is used; (b) oversampling via data augmentation (e.g., multi-scale training [35] and multi-scale inference [10, 34, 13]), while helpful, also introduces significant computational and storage overhead; (c) introducing scale-invariance priors (e.g., feature pyramids [21, 23, 22]), but this only partially alleviates scale bias, as both things and stuff vary widely in scale and image locations.
We propose a new method to address scale bias, orthogonal to previous techniques, using adaptive attentional processing. We warp images in-place while training models to shift the unbalanced scale distribution of objects (see Figure 2). While non-uniform warping techniques, i.e. allocating more pixels using a predefined guidance, have been applied to improve image classification [25], detection [36, 9, 37, 7] and segmentation [37] in supervised settings, they do not principally examine and address the scale bias for domain adaptation.
We propose Instance-level Saliency Guidance which analytically shifts the object scale distribution to remove bias or match target distribution by leveraging ground truth labels for maximum performance gain. Because the warping is in-place, we directly re-sample objects (e.g. expand smaller ones) and squeeze the large empty regions in the scene (e.g. road), allowing for a better use of the image pixel budget for balanced distribution training. We show that this warping improves backbone features of our model crucial for domain adaptation. To make sure our outputs are in the original unwarped configuration, we unwarp the features, as in LZU [37], and train without warping the object labels. During inference, no warping is performed and the method requires no additional computational latency. Our contributions are highlighted as follows:
We are the first work to show that scale bias in real-to-real domain adaptation can be effectively addressed through image warping, and that shifting source scale distribution significantly improves backbone features.
Our approach is agnostic to task (detection or segmentation), adaptation strategy, model architecture and warping guidance used for sampling.
Our approach improves domain adaptation across various lighting (+3.7 mAP50 for BDD100K Day Night detection), weather conditions (+6.1 mAP50 for BDD100K Clear DENSE Foggy detection), and geographies (+6.3 mIoU for Cityscapes ACDC semantic segmentation). See Figure 1.
Our approach is efficient: we introduce train-time image warping and feature unwarping for domain adaptation without need for test-time warping. This results in minimal training overhead and no additional inference latency.
2 Related Work
Alleviating Scale Bias. Recognition across different scales is a historically important problem. Priors that encourage scale invariance like feature (or anchor) pyramids are widely used in pre-deep-learning approaches [23, 5] and recently [27, 21, 22]. While helpful, these priors do not solve scale bias completely, as an object can be bigger or smaller than pre-defined scales. Common augmentation techniques such as multi-scale training [27, 35, 17, 13] can considerably increase training time. Similarly, test time scaling [10, 34] and multi-scale or multi-crop predictions [26, 13] improve accuracy at the cost of latency and memory [15]. Another line of works select “optimal” scale at inference [3, 8], but don’t improve the learned model itself. Compared to earlier work, our method explicitly addresses scale bias requiring (a) little additional memory during training since we warp images in-place; and (b) no additional inference latency as we do not warp during testing. Crucially, our approach is orthogonal to other scale bias reduction techniques, and can be used in a plug-and-play manner.
Guided Image Warping. Non-uniform Spatial Transformations [42] have been used to correct image distortions arising from camera acquisition. Recently, spatial transformation methods [25, 16] for learned models have advocated introducing image distortions instead. Learning to Zoom [25] showed over-sampling task-specific salient regions is useful as visual recognition requires combining information from different spatial resolutions. Building on this paradigm, warping for other tasks like detection [36, 37] and segmentation [37] were explored. Fovea [36], and Two Plane Prior [9] designed “salient” priors to guide warping, introducing static & temporal, and geometric saliency guidance respectively.
Our approach stands outs in the following ways: (a) Unlike previous image warping efforts focusing on supervised settings, our innovative training strategy and saliency guidance tailor image warping for unsupervised domain adaptation to directly address scale bias. (b) Unlike former image warping methods that indiscriminately process objects and background elements, we introduces an analytical instance-level saliency guidance that targets regions containing objects.
Image Transformations for Domain Adaptation. Image Transformations like flipping, cropping, blurring, color jittering, grayscaling, contrast adjustment are widely used as perturbations for domain adaptation [39, 45, 46, 24, 31, 48, 38, 6, 41, 43, 18, 12, 13, 11, 14, 17] to increase source domain diversity and decrease domain gap to the target domain. For example, 2PCNet [17] introduces a domain-specific Night Augmentation (NightAug) pipeline that attempts to make day images appear closer to night images. To the best of our knowledge, in-place image warping as a domain adaptive perturbation has not been explored in prior works. Our approach is agnostic to the specific domain adaptation training algorithm, and can be easily incorporated with any other augmentation strategy.
3 Method
We first describe how we warp images assuming a spatial guidance function while remaining task agnostic. Then, we design our warping guidance that can explicitly address scale bias (as shown in Figure 1). Finally, we detail how to train models with warped images that adapt better to new domains. The entire strategy is illustrated in Figure 3.
3.1 Image Warping and Feature Unwarping
We employ the saliency guided sampling mechanism from [25]. However, other methods may also be used (like thin plate spline transformations [16, 7]).
Image Warping: We consider an inverse transformation parameterized by saliency such that,
a resampled for every location is obtained by using warp given input image and saliency guidance , which is a 2D map of the output image size. The rationale and choice of is described in Section 3.2.
Feature Unwarping: If we operate in the space of warped images, our outputs would also be warped. However, we do not wish to warp our labels to match the input warped image, hence, we perform feature unwarping of spatial backbone features and train the model without warping our labels or unwarping our predictions, like [36, 9]. The benefit of this strategy is that it renders our warping mechanism task agnostic. We adopt the strategy from LZU [37], which approximates with , a piecewise tiling of invertible bilinear maps, and then perform feature unwarping via,
We restrict the space of allowable warps for such that axis alignment is preserved. Entire rows or columns are “stretched” or “compressed” in this case, which has been observed to perform better [36, 9, 37]:
where is a Gaussian kernel and denote the two image axes in the pixel space. and are marginalized versions of along and respectively.
3.2 Designing Saliency Guidance
Warping guidance decides which image regions are warped (e.g., scaled) more than others. Since images are warped in-place, if some regions are expanded, others must be reduced in size (Figure 4). For the warping mechanism considered [25, 36, 37], this requires designing a saliency function . We investigate three saliency functions, based on two approximate guidances proposed in previous work [36, 37, 9] and one that we propose in this work.
Approximate Saliency Guidance: The Geometric Prior [9] assumes that a dominant vanishing point exists in the distant scene, and designs a saliency map to exploit this property and oversample objects that appear small. But this prior tries to warp the entire scene that is farther away and the objects themselves may not be oversampled sufficiently to overcome size bias (Figure 4).
The Static Prior [36, 37] assumes that training set statistics will remain valid at test time, which is often not the case in domain adaptation due to domain shift and label unavailability. This incorrect assumption leads to the creation of a saliency map targeting average object locations, which may not accurately represent specific target images where objects deviate from these averages (Figure 4). Moreover, warping during the test phase, after adaptation, increases inference latency and is less effective due to the unknown scale distribution of the target domain (See Section 5.4 for ablation). Essentially, these saliency approaches do not utilize actual object labels and fail to concentrate on oversampling the objects, a strategy that would significantly benefit the shifting of the object scale distribution.
Instance-level Saliency Guidance. We propose Instance-level Saliency Guidance which explicitly shifts the object scale distribution by leveraging ground truth labels (Figure 4). During training, source image has corresponding labels available. For simplicity, we assume the labels are bounding boxes.
To shift the object scale distribution, we oversample image regions corresponding to every object instance. Following [1, 36], we use kernel density estimation to model the bounding box saliency as a sum of Gaussians,
where , and represent the center, width, and height of the bounding box, is the normal distribution and is the saliency scale or bandwidth.
We propose an expansion factor to decide the warping extent of all objects,
where is the fraction of object size corresponding to an interval between size thresholds and (See Figure 1). In our case, we consider , to represent the percentage of the small, medium, and large objects in the dataset, defined by COCO [20] thresholds for object sizes in the image.
Saliency scale and expansion factor are inversely proportional, thus where is a constant. We clip between of 0 and , since negative saliency values lead to holes, and an upper bound to discourage extreme warping.
When segmentation masks are available but not bounding boxes, we convert them to boxes, termed “from-seg” boxes. This conversion minimally impacts performance, as detailed in our ablations in Section 5.4 and Supplementary.
3.3 Warping for Domain Adaptation
We consider unsupervised domain adaptation, wherein source domain is available with labels, i.e. dataset and target domain only consists of images, .
Domain Adaption Framework: The standard framework comprises of two stages. First, a supervised pre-training stage where a model is trained using supervised loss on source domain, Second, in the unsupervised adaptation stage the model is adapted to target domain using dataset through an additional unsupervised loss. Generally, a student and teacher (initialized using ) are trained jointly through self-training, by employing psuedo labels [17, 12]. While training the student, a supervised loss is used to preserve performance on the source domain in a supervised branch:
The teacher is updated as the exponential moving average (EMA) of the student.
Warping and Unwarping in the Domain Adaptation Framework: The design decisions that guide our choices are: (a) Maximize transferability of features from supervised stage to unsupervised training stages. (b) Warping at test time after adaptation is likely to be suboptimal as the target domain’s scale distribution is unknown and can introduce latency. If the target distribution is known, the Instance-level guidance can be shifted analytically to match this distribution. Thus, our warping mechanism operates only where images from source domain are used, i.e. it is employed in the supervised pre-training stage and the student model in the unsupervised stage when source image is forwarded. We do experiment by warping target images too while employing Static [36, 37] and Geometric prior [9] saliency guidances (which do not use ground truth object labels to generate saliency), however, that reduces performance (See Table 8).
Consider the recognition model as a combination of backbone and prediction heads , i.e. . We warp the image and then unwarp the feature maps from backbone respectively to compose,
following [37] for a source image . Thus, losses are modified for the supervised pre-training stage as,
and for the unsupervised training stage as,
For our experiments, we use DAFormer [12] and 2PCNet [17] as our baseline segmentation and detection domain adaption algorithms respectively. We do not make additional assumptions about the domain adaptation setup, and thus our method is agnostic to the task specification (segmentation or detection). Other domain adaption algorithms [13, 14, 18, 6, 48] are also applicable.
Inference: We employ the teacher model during inference, and do not require any image warping and feature unwarping at test time. This is in contrast with image warping techniques proposed in the supervised setting [36, 9, 37], which incur overheads from test-time warping and assume additional information (like vanishing points [9]) to generate saliency maps at test time.
Supervised Setting: Our training method can also be applied in supervised learning scenario by following the supervised pre-training steps to train a model . That is, omitting the unsupervised training stage completely and only performing supervised pre-training stage – then removing the warping and unwarping layers completely from the trained model. This is unlike prior work [25, 36, 8, 37] which advocates warping at test time too. Model performance improves significantly on target domain without performing adaptation (See Table 5). For Source Domain Results, see Supplementary Section C.
4 Experimental Methodology and Datasets
We implement our source domain warping technique alongside a baseline domain adaptation algorithm, ensuring fair comparisons by using the same model, datasets, training schedules, hyperparameters, and random seeds.
Object Detection: Following the recent and popular domain adaptation strategy 2PCNet [17] (2023), we use Faster R-CNN [27] with ResNet-50 [10] adhering to their training hyperparameters and protocols. While 2PCNet [17] solely focuses on Day-to-Night adaptation, our method spans various adaptation scenarios. For scenarios other than day-to-night adaptation, we exclude the NightAug augmentation proposed in [17] for both the baseline and our method.
Semantic Segmentation: Aligning with the recent and popular domain adaptation strategy DAFormer [12] (2022), we employ the same SegFormer [44] head and MiT-B5 [44] backbone, and follow their training hyperparameters, protocols, and seed for fair comparison. While DAFormer [12] shows results on Sim2Real Gap which is not a focus of our work, results and discussion are available in Supplementary Section A. Extensions of DAFormer, including HRDA [13] and MIC [14], also benefits from our methodology, detailed in the Appendix B
4.1 Datasets
BDD100K [47] features 100,000 images with a resolution of 1280x720 for object detection and segmentation, covering various weather conditions and times of day. It includes 10 categories for annotation.
Cityscapes [4] provides 5,000 images of urban road scenes at a resolution of 2048x1024 with clear weather for semantic segmentation, with 19 categories.
DENSE [2] provides 12,997 images at a resolution of 1920x1024, capturing diverse weather conditions such as heavy fog and heavy snow.
ACDC [32] is designed for adverse conditions (fog and snow), including 1,600 images at 2048x1024 for segmentation across 19 categories.
DarkZurich [30] is tailored for low-light conditions, offering 2,416 unlabeled nighttime images and 151 labeled twilight images for segmentation, with a resolution of 1920x1080, and a focus on urban settings.
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl | person | car | truck | bike |
2PCNet | 27.0 | 45.0 | 30.1 | 0.0 | 15.4 | 35.5 | 87.5 | 72.5 | 19.1 | 0.9 |
+ Ours w/ Sta. Prior | 28.6 | 48.5 | 31.7 | 0.0 | 19.6 | 34.0 | 88.1 | 72.5 | 25.6 | 7.7 |
+ Ours w/ Geo. Prior | 27.1 | 45.8 | 29.5 | 0.0 | 15.2 | 37.5 | 86.5 | 75.4 | 21.0 | 0.4 |
+ Ours w/ Inst. Warp | 33.1 | 51.1 | 39.7 | 0.0 | 14.5 | 45.2 | 85.8 | 66.3 | 50.0 | 2.2 |
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl | person | rider | car | truck | bus |
bike |
2PCNet | 25.8 | 48.4 | 24.4 | 10.9 | 34.1 | 49.0 | 56.3 | 40.4 | 77.0 | 57.2 | 55.9 | 36.8 | 36.8 | 55.6 | 68.1 | ||||||
+ Ours w/ Sta. Prior | 26.8 | 50.6 | 24.9 | 11.4 | 35.9 | 49.6 | 57.4 | 42.5 | 77.4 | 60.1 | 59.5 | 44.8 | 39.4 | 55.7 | 68.9 | ||||||
+ Ours w/ Geo. Prior | 26.9 | 49.9 | 24.5 | 11.8 | 35.6 | 50.1 | 56.1 | 44.6 | 77.5 | 58.0 | 58.6 | 33.3 | 45.8 | 55.8 | 69.5 | ||||||
+ Ours w/ Inst. Warp | 27.1 | 51.4 | 24.6 | 11.0 | 36.7 | 49.0 | 57.2 | 46.6 | 77.5 | 59.8 | 60.9 | 46.3 | 42.2 | 54.9 | 68.3 |
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl | person | rider | car | truck | bus |
bike |
2PCNet | 23.5 | 46.4 | 21.1 | 9.5 | 26.0 | 43.3 | 54.4 | 30.8 | 73.1 | 53.8 | 55.2 | 37.5 | 44.5 | 49.4 | 65.2 | ||||||
+ Ours w/ Sta. Prior | 25.3 | 49.4 | 22.4 | 10.3 | 27.4 | 45.1 | 58.2 | 39.3 | 75.0 | 53.6 | 54.5 | 41.4 | 49.6 | 52.9 | 69.6 | ||||||
+ Ours w/ Geo. Prior | 25.6 | 49.7 | 22.9 | 10.0 | 27.7 | 45.7 | 58.8 | 41.8 | 75.0 | 55.8 | 55.2 | 40.8 | 49.3 | 50.5 | 69.4 | ||||||
+ Ours w/ Inst. Warp | 25.6 | 50.1 | 22.6 | 9.8 | 27.6 | 46.6 | 58.9 | 42.7 | 75.1 | 55.0 | 55.5 | 42.6 | 50.1 | 51.9 | 69.0 |
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl | person | rider | car | truck | bus |
bike | ||
2PCNet | 16.6 | 30.8 | 14.9 | 6.0 | 24.8 | 26.6 | 41.0 | 20.0 | 77.0 | 41.8 | 17.1 | 29.9 | 19.3 | ||
+ Ours w/ Sta. Prior | 17.3 | 31.0 | 16.7 | 6.2 | 25.9 | 29.5 | 42.2 | 20.9 | 76.7 | 40.5 | 21.7 | 24.4 | 21.8 | ||
+ Ours w/ Geo. Prior | 16.6 | 30.3 | 15.4 | 5.2 | 25.1 | 28.7 | 42.3 | 19.2 | 75.0 | 39.5 | 18.1 | 26.7 | 21.3 | ||
+ Ours w/ Inst. Warp | 17.9 | 32.1 | 18.2 | 6.8 | 26.7 | 32.1 | 45.7 | 20.6 | 77.6 | 43.2 | 20.8 | 26.7 | 22.2 |
5 Experiments
Method | mIoU | road |
building | wall | fence | pole |
terrain | sky | person | rider | car | truck | bus | train |
bike | ||||||||||
ACDC Val Split: After Supervised Pre-Training | ||||||||||||||||||||||||||||||
DAFormer [12] | 55.8 | 77.6 | 57.4 | 77.9 | 39.4 | 30.4 | 48.3 | 54.5 | 46.4 | 70.0 | 37.9 | 72.2 | 52.6 | 32.0 | 81.4 | 72.7 | 73.7 | 70.4 | 37.6 | 27.3 | ||||||||||
+ Ours w/ Sta. | 55.8 | 77.5 | 55.8 | 76.7 | 39.8 | 28.7 | 48.6 | 56.0 | 43.9 | 69.0 | 37.5 | 73.0 | 53.6 | 32.9 | 81.1 | 71.8 | 72.6 | 82.5 | 36.0 | 22.8 | ||||||||||
+ Ours w/ Geo. | 55.7 | 80.4 | 57.3 | 78.3 | 37.7 | 28.4 | 48.2 | 54.4 | 44.8 | 70.3 | 37.4 | 75.1 | 52.8 | 30.5 | 80.7 | 75.6 | 72.8 | 65.8 | 38.9 | 29.8 | ||||||||||
+ Ours w/ Inst. | 56.8 | 80.2 | 48.3 | 76.0 | 35.0 | 31.6 | 48.6 | 58.2 | 45.3 | 69.4 | 36.9 | 77.2 | 53.2 | 26.2 | 82.0 | 77.6 | 80.0 | 69.4 | 39.7 | 43.7 | ||||||||||
ACDC Val Split: After Unsupervised Domain Adaptation | ||||||||||||||||||||||||||||||
DAFormer [12] | 57.6 | 72.7 | 57.5 | 80.1 | 42.5 | 38.0 | 50.9 | 45.1 | 50.0 | 71.1 | 38.5 | 67.0 | 56.0 | 29.9 | 81.8 | 76.6 | 78.9 | 79.9 | 40.8 | 36.7 | ||||||||||
+ Ours w/ Sta. | 59.1 | 73.8 | 56.8 | 79.1 | 45.2 | 39.6 | 50.0 | 48.3 | 49.4 | 71.0 | 39.3 | 68.6 | 55.8 | 36.9 | 82.1 | 73.2 | 81.3 | 88.7 | 37.1 | 47.4 | ||||||||||
+ Ours w/ Geo. | 58.4 | 76.4 | 57.6 | 79.9 | 42.2 | 39.0 | 51.7 | 46.4 | 49.5 | 72.0 | 39.0 | 71.1 | 56.5 | 30.4 | 82.7 | 79.3 | 79.1 | 69.8 | 41.5 | 44.7 | ||||||||||
+ Ours w/ Inst. | 61.8 | 82.9 | 56.1 | 79.8 | 44.6 | 40.3 | 52.7 | 60.8 | 52.5 | 72.0 | 38.4 | 78.0 | 56.6 | 30.5 | 84.9 | 80.2 | 86.9 | 86.4 | 44.5 | 45.8 | ||||||||||
ACDC Test Split Results | ||||||||||||||||||||||||||||||
ADVENT [40] | 32.7 | 72.9 | 14.3 | 40.5 | 16.6 | 21.2 | 9.3 | 17.4 | 21.2 | 63.8 | 23.8 | 18.3 | 32.6 | 19.5 | 69.5 | 36.2 | 34.5 | 46.2 | 26.9 | 36.1 | ||||||||||
MGCDA [30] | 48.7 | 73.4 | 28.7 | 69.9 | 19.3 | 26.3 | 36.8 | 53.0 | 53.3 | 75.4 | 32.0 | 84.6 | 51.0 | 26.1 | 77.6 | 43.2 | 45.9 | 53.9 | 32.7 | 41.5 | ||||||||||
DANNet [43] | 50.0 | 84.3 | 54.2 | 77.6 | 38.0 | 30.0 | 18.9 | 41.6 | 35.2 | 71.3 | 39.4 | 86.6 | 48.7 | 29.2 | 76.2 | 41.6 | 43.0 | 58.6 | 32.6 | 43.9 | ||||||||||
DAFormer [12] | 55.4 | 58.4 | 51.3 | 84.0 | 42.7 | 35.1 | 50.7 | 30.0 | 57.0 | 74.8 | 52.8 | 51.3 | 58.3 | 32.6 | 82.7 | 58.3 | 54.9 | 82.4 | 44.1 | 50.7 | ||||||||||
+ Ours w/ Inst. | 61.7 | 83.0 | 53.2 | 85.5 | 47.4 | 38.3 | 46.0 | 51.4 | 57.8 | 73.9 | 56.2 | 82.1 | 61.3 | 32.3 | 85.5 | 69.0 | 68.9 | 82.5 | 46.7 | 52.0 |
Method | mIoU | road |
building | wall | fence | pole |
terrain | sky | person | rider | car | truck | bus | train |
bike | ||||||||||
DarkZurich Val Split: After Supervised Pre-Training | ||||||||||||||||||||||||||||||
DAFormer | 30.0 | 90.7 | 52.7 | 45.5 | 18.4 | 39.0 | 38.4 | 27.5 | 18.3 | 59.0 | 26.5 | 11.1 | 20.4 | 21.1 | 58.6 | - | - | - | 22.1 | 20.0 | ||||||||||
+ Ours w/ Sta. | 32.7 | 91.8 | 57.2 | 47.3 | 19.3 | 46.2 | 40.1 | 21.0 | 18.2 | 53.5 | 31.9 | 6.4 | 20.6 | 18.6 | 64.1 | - | - | - | 25.0 | 26.4 | ||||||||||
+ Ours w/ Geo. | 30.2 | 89.9 | 50.5 | 45.9 | 20.5 | 36.8 | 39.1 | 21.0 | 19.8 | 54.7 | 28.4 | 7.9 | 19.0 | 31.1 | 58.4 | - | - | - | 18.1 | 33.4 | ||||||||||
+ Ours w/ Inst. | 34.0 | 91.9 | 59.4 | 49.3 | 27.7 | 39.2 | 39.4 | 27.6 | 15.8 | 55.1 | 29.9 | 11.4 | 20.7 | 9.2 | 61.9 | - | - | - | 27.9 | 44.7 | ||||||||||
DarkZurich Val Split: After Unsupervised Domain Adaptation | ||||||||||||||||||||||||||||||
DAFormer | 33.8 | 92.0 | 66.7 | 47.3 | 25.9 | 50.8 | 38.0 | 24.6 | 19.4 | 62.2 | 31.9 | 17.9 | 20.4 | 28.4 | 61.7 | - | - | - | 21.0 | 34.3 | ||||||||||
+ Ours w/ Sta. | 33.8 | 93.3 | 68.5 | 50.0 | 28.4 | 54.3 | 41.7 | 19.9 | 22.7 | 54.0 | 41.5 | 14.1 | 22.9 | 7.2 | 62.6 | - | - | - | 24.6 | 36.8 | ||||||||||
+ Ours w/ Geo. | 33.3 | 84.3 | 60.3 | 60.2 | 23.9 | 49.4 | 38.7 | 12.7 | 25.0 | 52.0 | 39.5 | 20.8 | 37.1 | 16.2 | 64.6 | - | - | - | 21.1 | 27.8 | ||||||||||
+ Ours w/ Inst. | 37.1 | 88.7 | 70.9 | 60.1 | 42.1 | 49.6 | 39.8 | 47.9 | 21.5 | 49.8 | 35.7 | 25.4 | 23.8 | 25.9 | 66.8 | - | - | - | 24.3 | 32.4 | ||||||||||
DarkZurich Test Split Results | ||||||||||||||||||||||||||||||
ADVENT [40] | 29.7 | 85.8 | 37.9 | 55.5 | 27.7 | 14.5 | 23.1 | 14.0 | 21.1 | 32.1 | 8.7 | 2.0 | 39.9 | 16.6 | 64.0 | 13.8 | 0.0 | 58.8 | 28.5 | 20.7 | ||||||||||
MGCDA [30] | 42.5 | 80.3 | 49.3 | 66.2 | 7.8 | 11.0 | 41.4 | 38.9 | 39.0 | 64.1 | 18.0 | 55.8 | 52.1 | 53.5 | 74.7 | 66.0 | 0.0 | 37.5 | 29.1 | 22.7 | ||||||||||
DANNet [43] | 44.3 | 90.0 | 54.0 | 74.8 | 41.0 | 21.1 | 25.0 | 26.8 | 30.2 | 72.0 | 26.2 | 84.0 | 47.0 | 33.9 | 68.2 | 19.0 | 0.3 | 66.4 | 38.3 | 23.6 | ||||||||||
DAFormer [12] | 53.8 | 93.5 | 65.5 | 73.3 | 39.4 | 19.2 | 53.3 | 44.1 | 44.0 | 59.5 | 34.5 | 66.6 | 53.4 | 52.7 | 82.1 | 52.7 | 9.5 | 89.3 | 50.5 | 38.5 | ||||||||||
+ Ours w/ Inst. | 54.4 | 94.2 | 68.1 | 72.6 | 43.2 | 23.6 | 51.0 | 48.1 | 48.5 | 60.6 | 33.6 | 54.8 | 55.5 | 49.0 | 79.8 | 51.1 | 13.2 | 90.2 | 53.2 | 43.4 |
5.1 Domain Adaptive Object Detection
Tables 1, 2, 3 and 4 show improvements of our method over state-of-the-art 2PCNet [17] across lighting, weather and geographic adaptations respectively. Moroever, our instance-level saliency guidance always improves over approximate saliency guidance [36, 9]. Note that 2PCNet [17] focused solely on day night adaptation with it’s NightAug augmentation, which was only used in the BDD100K day night experiment (Table 3) and omitted for others. Our method to address scale bias is more general in comparison.
5.2 Domain Adaptive Semantic Segmentation
To demonstrate our method is task agnostic, we incorporate it within domain adaptive semantic segmentation using the popular DAFormer [12] algorithm. The resulting approach shows significant improvement on both Val and Test Splits of the ACDC dataset (Table 5). Here, saliency guidance choice matters: our instance-level image saliency guidance outperforms Static Prior [36] and Geometric Prior [9]. Being agnostic to the DA algorithm, our method also boosts DAFormer extensions like HRDA [13] and MIC [14] (See Supplementary Section B). These extensions mitigate scale bias to some extent but add significant training memory (+5GB) and inference latency (+800 ms on RTX 4090 GPU), whereas our approach requires negligible additional training memory (+0.04 GB) and adds no additional inference latency.
Table 5 also shows our method’s superiority across scenarios, with or without unsupervised domain adaptation. This suggests that the learned backhone features due to our method generalize better to unseen domains. Additionally, Figure 7 illustrates superior performance of our method in identifying both background elements (e.g., sky, sidewalks, buildings) and foreground objects (e.g., traffic lights and traffic signs).
5.3 Visual Diagnosis of Learned Models
In Section 5.2, we noted that the learned backbone feature of our method generalize better. To corroborate this claim, we employed GradCAM [33] visualizations, showcasing ResNet features from detectors trained on BDD100K [47] in Figure 8. The heatmap for models trained with 2PCNet without warping reveals a dispersed focus across the image, indicating a lack of precision. The choice of saliency guidance plays a critical role: warping guided by Geometric Prior [9] results in less focus on the target (e.g., cars) and unnecessary background emphasis. Conversely, our instance-level saliency guidance achieves a focused heatmap on relevant objects, minimizing irrelevant activations.
5.4 Ablation Study
Parameters | aAcc | mIoU | mAcc |
P=, U=1.00 | 84.7 | 60.1 | 73.2 |
P=, U=1.00 | 84.2 | 61.4 | 74.1 |
P=, U=0.50 | 85.2 | 60.1 | 72.5 |
P=, U=0.75 | 82.4 | 60.3 | 73.1 |
P=, U=1.00 | 85.3 | 61.8 | 73.8 |
Sample | Percent | Method | aAcc | mIoU | mAcc |
Uniform | 25% | SegFormer | 95.0 | 71.0 | 79.1 |
Ours | 95.1 | 71.9 | 80.0 | ||
50% | SegFormer | 95.3 | 73.6 | 81.4 | |
Ours | 95.5 | 74.4 | 82.1 | ||
None | 100% | SegFormer | 95.6 | 75.3 | 82.9 |
Ours | 95.9 | 76.8 | 84.2 |
For all ablation studies, we utilize the Cityscape ACDC domain adaptive semantic segmentation task.
Saliency Product and Upperbound . Varying and shows optimal performance at and , according to quantitative scores in Table 7(a). Hence, and are adopted for all major experiments.
Warping on Source vs. Target. Given the necessity of ground truth labels for our instance-level saliency guidance, warping in the target domain for unsupervised domain adaptation is infeasible. Nevertheless, as shown in Table 8 warping source domain images with our saliency guidance surpasses performance using approximated saliency [36, 9] with both source and target domain warping.
Warping with From-Seg vs. Groundtruth bboxes. For object detection, we utilize ground truth bounding boxes for saliency guidance. However, semantic segmentation tasks often lack this data, leading us to use bounding boxes computed from segmentations, “from-seg” for short. Results indicate superior performance with ground truth boxes, but “from-seg” boxes are a close second, highlighting their effectiveness when ground truth is unavailable (Table 8).
Distribution Balancing by Undersampling. We analyzed undersampling strategies that balance the percentages of small, medium, and large objects, aiming to create a more uniform distribution that mitigates scale bias. As shown in Table 7(b), our experiments applying 25% and 50% of undersampling to the Cityscapes dataset indicates that models trained on these “uniform” subsets perform worse than those trained on the complete dataset, highlighting the challenges of addressing scale bias through simple subsampling.
Method | Warp | Bboxes | aAcc | mIoU | mAcc |
DAFormer | - | - | 81.9 | 57.6 | 70.6 |
+ Ours w/ Sta. Prior | src & tgt & test | - | 82.5 | 55.3 | 68.3 |
src & tgt | - | 83.8 | 57.4 | 69.6 | |
src | - | 82.3 | 59.1 | 73.3 | |
+ Ours w/ Geo. Prior | src & tgt & test | - | 82.2 | 57.7 | 71.6 |
src & tgt | - | 81.9 | 56.5 | 70.9 | |
src | - | 83.2 | 58.4 | 70.9 | |
+ Ours w/ Inst. (tgt & test inapplicable) | src | from-seg | 84.0 | 60.1 | 73.5 |
src | gt | 85.3 | 61.8 | 73.8 |
6 Limitations and Conclusion
In this work, we address scale bias with attentional processing, using saliency-guided image warping that integrates seamlessly into any unsupervised domain adaptation method. Leveraging scale statistics from the source dataset and ground truth labels, our approach effectively navigates challenges in lighting, weather, and geographic differences. Extensive experiments demonstrate the effectiveness of our warping method and instance-level saliency guidance. Importantly, our approach adds no inference latency. Depending on the available information (e.g., vanishing point, target statistics), other priors could be used with our approach, showing improvement over baselines, albeit to a lesser extent.
One limitation of our method arises in densely populated scenes (e.g., New York’s Times Square), where the limited background space challenges the expansion and contraction capabilities of our warping. Besides, our approach is less effective on certain synthetic datasets, such as Synthia [29], which do not exhibit realistic scale bias. However, our approach proves effective on other synthetic datasets like GTA5 [28], a role-playing game dataset featuring city driving components with realistic traffic and exhibiting realistic scale biases.
Societal Impact. While our method demonstrates improved perception on publicly available night and adverse weather driving datasets, extensive testing is required before real-world deployment on self-driving vehicles.
Acknowledgment. This work was supported in parts by a sponsored research contract from General Motors Research-Israel, NSF grant CNS-2038612, and DOT RITA Mobility-21 Grant 69A3551747111.
Appendix A Synthetic to Real Domain Adaptation
Our proposed method addresses source scale biases that are present in natural scene datasets. Synthetic To Real (Sim2Real) Domain Gap is not a focus of our work as a synthetic scene dataset may or may not exhibit the same bias.
However, we do explore if our method improves performance on the Sim2Real Domain Gap by considering GTA [28] and Synthia [29] synthetic datasets. Table 9 and Table 10 show the results for GTA Cityscapes and Synthia Cityscapes domain adaptation respectively.
Our method improves GTA Cityscapes adaptation considerably, while improvement on Synthia Cityscapes adaptation is marginal, which corresponds to our intuition regarding both the source datasets (See Fig 9).
Grand Theft Auto 5 (GTA) is a role playing game with a city driving component along with realistic traffic, thus GTA images are close to real driving scenes and hence exhibits similar scale biases. Synthia, on the other hand, is a rendered dataset of a virtual city, and does not have realistic traffic simulation of dynamic objects like cars and people. Morever, many images are not captured from a vehicle’s perspective. Thus, images are not close to real driving scenes and do not exhibit scale bias.
Method | mIoU | road |
building | wall | fence | pole |
terrain | sky | person | rider | car | truck | bus | train |
bike | ||||||||||
CBST [50] | 45.9 | 91.8 | 53.5 | 80.5 | 32.7 | 21.0 | 34.0 | 28.9 | 20.4 | 83.9 | 34.2 | 80.9 | 53.1 | 24.0 | 82.7 | 30.3 | 35.9 | 16.0 | 25.9 | 42.8 | ||||||||||
DACS [38] | 52.1 | 89.9 | 39.7 | 87.9 | 30.7 | 39.5 | 38.5 | 46.4 | 52.8 | 88.0 | 44.0 | 88.8 | 67.2 | 35.8 | 84.5 | 45.7 | 50.2 | 0.0 | 27.3 | 34.0 | ||||||||||
CorDA [41] | 56.6 | 94.7 | 63.1 | 87.6 | 30.7 | 40.6 | 40.2 | 47.8 | 51.6 | 87.6 | 47.0 | 89.7 | 66.7 | 35.9 | 90.2 | 48.9 | 57.5 | 0.0 | 39.8 | 56.0 | ||||||||||
ProDA [48] | 57.5 | 87.8 | 56.0 | 79.7 | 46.3 | 44.8 | 45.6 | 53.5 | 53.5 | 88.6 | 45.2 | 82.1 | 70.7 | 39.2 | 88.8 | 45.5 | 59.4 | 1.0 | 48.9 | 56.4 | ||||||||||
DAFormer [12] | 68.3 | 95.7 | 70.2 | 89.4 | 53.5 | 48.1 | 49.6 | 55.8 | 59.4 | 89.9 | 47.9 | 92.5 | 72.2 | 44.7 | 92.3 | 74.5 | 78.2 | 65.1 | 55.9 | 61.8 | ||||||||||
DAFormer (seed=0) | 66.9 | 92.6 | 58.9 | 89.3 | 54.2 | 42.7 | 49.4 | 57.0 | 55.8 | 89.2 | 49.8 | 89.5 | 72.7 | 41.7 | 92.0 | 62.0 | 82.8 | 71.3 | 56.5 | 62.9 | ||||||||||
+ Ours | 68.5 | 92.9 | 60.0 | 89.8 | 55.9 | 51.5 | 49.0 | 57.2 | 62.2 | 89.6 | 50.2 | 91.5 | 71.9 | 44.8 | 93.0 | 78.7 | 79.8 | 63.6 | 56.6 | 63.6 |
Method | mIoU | road |
building | wall | fence | pole |
terrain | sky | person | rider | car | truck | bus | train |
bike | ||||||||||
CBST [50] | 42.6 | 68.0 | 29.9 | 76.3 | 10.8 | 1.4 | 33.9 | 22.8 | 29.5 | 77.6 | - | 78.3 | 60.6 | 28.3 | 81.6 | - | 23.5 | - | 18.8 | 39.8 | ||||||||||
DACS [38] | 48.3 | 80.6 | 25.1 | 81.9 | 21.5 | 2.9 | 37.2 | 22.7 | 24.0 | 83.7 | - | 90.8 | 67.6 | 38.3 | 82.9 | - | 38.9 | - | 28.5 | 47.6 | ||||||||||
CorDA [41] | 55.0 | 93.3 | 61.6 | 85.3 | 19.6 | 5.1 | 37.8 | 36.6 | 42.8 | 84.9 | - | 90.4 | 69.7 | 41.8 | 85.6 | - | 38.4 | - | 32.6 | 53.9 | ||||||||||
ProDA [48] | 55.5 | 87.8 | 45.7 | 84.6 | 37.1 | 0.6 | 44.0 | 54.6 | 37.0 | 88.1 | - | 84.4 | 74.2 | 24.3 | 88.2 | - | 51.1 | - | 40.5 | 45.6 | ||||||||||
DAFormer [12] | 60.9 | 84.5 | 40.7 | 88.4 | 41.5 | 6.5 | 50.0 | 55.0 | 54.6 | 86.0 | - | 89.8 | 73.2 | 48.2 | 87.2 | - | 53.2 | - | 53.9 | 61.7 | ||||||||||
DAFormer (seed=0) | 59.2 | 87.8 | 47.6 | 87.6 | 43.7 | 5.8 | 49.0 | 48.0 | 53.1 | 82.3 | - | 71.8 | 71.7 | 46.0 | 87.5 | - | 49.3 | - | 52.7 | 63.4 | ||||||||||
+ Ours | 59.4 | 92.7 | 57.7 | 87.1 | 44.8 | 7.1 | 46.2 | 46.9 | 53.6 | 80.6 | - | 68.2 | 72.5 | 42.5 | 90.6 | - | 41.6 | - | 55.2 | 63.3 |
Appendix B Comparison with other Domain Adaptation Strategies
Method | Training Memory | Inference Time |
DAFormer | 16.41 GB | 195.0 ms |
DAFormer + Ours | + 0.04 GB | + 0 ms |
DAFormer + HRDA | + 5.11 GB | + 801.2 ms |
Domain adaptation extensions have been proposed on the DAFormer [12] training framework, such as HRDA [13] and MIC [14] which also show good performance.
Differences from DAFormer. HRDA [13] proposed a multi-scale high-resolution crop training strategy along with sliding window inference. MIC [14] proposed a masking consistency strategy for target domain images to improve spatial context.
HRDA [13] departs from DAFormer [12] by increasing training and inference image scales by a factor of 2 and performing sliding window inference (a test time scale augmentation). Training and inference computational costs significantly increase with HRDA’s strategy [13]. Correspondingly, our inference costs are same as DAFormer [12] (as we do not warp during test-time) while training time and memory consumption is marginally higher (See Table 11). Lastly, instance-level image warping is orthogonal to the aforementioned training strategies and can be used simultaneously with other training strategies.
Evaluation Methodology. Based on the above discussion, we apply HRDA’s and MIC’s strategy on DAFormer, but for a fair comparison, we perform training and inference following DAFormer’s [12] training/testing image scales and evaluation paradigm. The alternative comparison would be to train and evaluate on full-scale images for parity with HRDA, however, we do not have a GPU with high memory availability (like NVIDIA Tesla A40 with 48 GB or Tesla A100 with 80 GB memory).
Results. Table 12 and 13 present results for Cityscapes DarkZurich and Cityscapes ACDC respectively. MIC (HRDA) corresponds to MIC and HRDA training strategies added to DAFormer. We compare with DAFormer and MIC (HRDA) with and without including our Instance-level image warping approach. We observe smaller improvements in conjunction with our method with MIC (HRDA) as HRDA uses multi-scale cropping strategies, which also attempts to alleviate scale bias to some extent.
As our instance-level image warping approach is orthogonal to other training strategies and can be used in plug-and-play manner, we studied the interaction of our method with MIC [14] and HRDA [13] individually (See Table 14).
We observed that our method when employed in conjunction with MIC and HRDA consistently improved performance, however, was not able to improve upon the scenario where our method is applied to DAFormer directly. Moreover, HRDA failed to improve upon DAFormer itself when trained on DAFormer’s training scale (1024 512) instead of full scale (2048 1024).
We concluded that high resolution training is necessary for HRDA as posited [13] and the strategy does not work at all image scales. Further, we observed HRDA performs two kinds of cropping – a HR-detail crop to better segment small object details and LR-context crop to adapt better to large objects. At DAFormer’s training scales, the cropped size of the detail crop would be very small and hurt performance. The advantage of our instance-level warping method is that it increases focus on small object details, thus we replaced detail crop with our instance level image warping. Removing detail crop from HRDA improved performance over DAFormer itself, and in conjunction with our method showed the best performance.
Method | mIoU | road |
building | wall | fence | pole |
terrain | sky | person | rider | car | truck | bus | train |
bike | ||||||||||
DAFormer | 33.8 | 92.0 | 66.7 | 47.3 | 25.9 | 50.8 | 38.0 | 24.6 | 19.4 | 62.2 | 31.9 | 17.9 | 20.4 | 28.4 | 61.7 | - | - | - | 21.0 | 34.3 | ||||||||||
DAFormer + Ours | 37.1 | 88.7 | 70.9 | 60.1 | 42.1 | 49.6 | 39.8 | 47.9 | 21.5 | 49.8 | 35.7 | 25.4 | 23.8 | 25.9 | 66.8 | - | - | - | 24.3 | 32.4 | ||||||||||
MIC (HRDA) | 39.8 | 78.8 | 13.0 | 83.1 | 46.7 | 52.2 | 42.2 | 44.5 | 28.8 | 64.3 | 35.4 | 82.8 | 24.9 | 33.8 | 62.8 | - | - | - | 25.9 | 37.3 | ||||||||||
MIC (HRDA) + Ours | 40.1 | 75.8 | 5.8 | 83.3 | 55.7 | 54.8 | 44.3 | 30.5 | 38.0 | 66.3 | 42.7 | 82.4 | 39.4 | 15.1 | 69.3 | - | - | - | 25.1 | 32.9 |
Method | mIoU | road |
building | wall | fence | pole |
terrain | sky | person | rider | car | truck | bus | train |
bike | ||||||||||
DAFormer | 57.6 | 72.7 | 57.5 | 80.1 | 42.5 | 38.0 | 50.9 | 45.1 | 50.0 | 71.1 | 38.5 | 67.0 | 56.0 | 29.9 | 81.8 | 76.6 | 78.9 | 79.9 | 40.8 | 36.7 | ||||||||||
DAFormer + Ours | 61.8 | 82.9 | 56.1 | 79.8 | 44.6 | 40.3 | 52.7 | 60.8 | 52.5 | 72.0 | 38.4 | 78.0 | 56.6 | 30.5 | 84.9 | 80.2 | 86.9 | 86.4 | 44.5 | 45.8 | ||||||||||
MIC (HRDA) | 59.6 | 66.4 | 55.8 | 84.8 | 54.6 | 41.3 | 49.7 | 41.2 | 53.6 | 78.3 | 37.7 | 67.4 | 58.3 | 24.0 | 83.6 | 70.1 | 88.8 | 87.9 | 40.7 | 48.4 | ||||||||||
MIC (HRDA) + Ours | 61.6 | 80.6 | 38.6 | 83.7 | 50.5 | 41.3 | 50.5 | 56.7 | 49.6 | 75.2 | 39.4 | 83.9 | 58.9 | 32.0 | 86.0 | 75.4 | 91.8 | 88.0 | 40.2 | 48.5 |
Method | mIoU | road |
building | wall | fence | pole |
terrain | sky | person | rider | car | truck | bus | train |
bike | ||||||||||
DAFormer | 57.6 | 72.7 | 57.5 | 80.1 | 42.5 | 38.0 | 50.9 | 45.1 | 50.0 | 71.1 | 38.5 | 67.0 | 56.0 | 29.9 | 81.8 | 76.6 | 78.9 | 79.9 | 40.8 | 36.7 | ||||||||||
DAFormer + Ours | 61.8 | 82.9 | 56.1 | 79.8 | 44.6 | 40.3 | 52.7 | 60.8 | 52.5 | 72.0 | 38.4 | 78.0 | 56.6 | 30.5 | 84.9 | 80.2 | 86.9 | 86.4 | 44.5 | 45.8 | ||||||||||
MIC (DAFormer) | 58.8 | 61.1 | 59.5 | 73.2 | 47.4 | 45.2 | 51.4 | 44.7 | 48.2 | 78.4 | 38.1 | 51.4 | 60.4 | 41.3 | 84.3 | 78.5 | 84.3 | 78.9 | 43.9 | 46.5 | ||||||||||
MIC (DAFormer) + Ours | 60.6 | 72.8 | 62.9 | 73.4 | 45.1 | 36.5 | 52.9 | 49.0 | 49.9 | 76.9 | 39.8 | 65.5 | 60.6 | 40.4 | 85.2 | 80.9 | 90.5 | 87.0 | 41.0 | 41.3 | ||||||||||
HRDA (DAFormer) | 56.9 | 79.9 | 37.8 | 81.1 | 45.6 | 33.9 | 47.8 | 47.3 | 47.1 | 74.3 | 37.1 | 84.0 | 47.7 | 17.6 | 84.2 | 69.1 | 88.2 | 75.1 | 37.6 | 45.1 | ||||||||||
HRDA (DAFormer) + Ours | 57.7 | 85.6 | 48.2 | 71.7 | 41.6 | 39.4 | 50.8 | 17.7 | 47.8 | 75.2 | 37.9 | 81.4 | 56.3 | 25.0 | 82.3 | 73.4 | 88.8 | 82.3 | 45.3 | 46.7 | ||||||||||
HRDA (DAFormer) | 58.3 | 68.5 | 59.4 | 82.8 | 50.4 | 40.8 | 50.3 | 42.4 | 44.4 | 77.6 | 38.0 | 69.4 | 55.7 | 27.2 | 83.2 | 77.4 | 78.2 | 79.0 | 34.9 | 48.4 | ||||||||||
HRDA (DAFormer) + Ours | 62.1 | 89.7 | 61.1 | 83.9 | 43.4 | 39.5 | 52.7 | 43.1 | 45.0 | 75.6 | 38.7 | 86.1 | 55.0 | 28.0 | 84.9 | 81.1 | 88.5 | 86.0 | 44.9 | 53.4 |
Appendix C Supervised Setting
In Section 3.2, we mentioned that source pre-training also improves performance in the supervised setting on the source domain itself. We show improvements in segmentation performance on Cityscapes, when applied to SegFormer model (base architecture of the DAFormer [12] training strategy), see Table 15. Please observe visual results of our method in Figure 10. We also show improvements in detection performance on BDD100K (Day) and BDD100 (Clear) when applied to Faster R-CNN (base architecture of 2PCNet [17] training strategy), see Table 16 and Table 17. BDD100K (Day) includes both clear and bad weather images taken during the day, while BDD100K (Clear) includes both day and night images taken in clear weather.
Method | mIoU | road |
building | wall | fence | pole |
terrain | sky | person | rider | car | truck | bus | train |
bike | ||||||||||
SegFormer | 75.3 | 98.0 | 83.4 | 91.8 | 59.6 | 59.6 | 57.5 | 64.0 | 74.4 | 91.9 | 62.8 | 94.6 | 77.6 | 56.0 | 93.7 | 81.6 | 81.4 | 70.1 | 59.2 | 73.3 | ||||||||||
+ Ours w/ Sta. Prior | 76.1 | 97.9 | 83.6 | 91.9 | 58.0 | 58.0 | 57.7 | 63.3 | 73.9 | 91.7 | 64.4 | 94.4 | 77.7 | 57.4 | 93.8 | 81.0 | 87.7 | 80.3 | 60.7 | 72.8 | ||||||||||
+ Ours w/ Geo. Prior | 76.5 | 98.0 | 84.3 | 92.2 | 61.4 | 57.6 | 59.3 | 64.7 | 74.0 | 91.9 | 65.3 | 94.7 | 78.2 | 55.3 | 93.9 | 83.1 | 86.5 | 81.1 | 60.1 | 72.1 | ||||||||||
+ Ours w/ Inst. | 76.8 | 98.1 | 84.8 | 92.2 | 59.9 | 58.3 | 59.6 | 65.1 | 75.4 | 92.3 | 66.2 | 94.8 | 78.2 | 55.3 | 94.2 | 82.0 | 85.7 | 81.3 | 61.8 | 74.4 |
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl | person | rider | car | truck | bus |
bike |
FRCNN | 30.1 | 56.4 | 28.1 | 13.9 | 37.6 | 51.0 | 64.0 | 50.7 | 80.3 | 62.5 | 62.9 | 45.3 | 49.7 | 66.7 | 69.8 | ||||||
+ Ours w/ Sta. Prior | 30.9 | 57.1 | 28.6 | 14.4 | 38.9 | 53.4 | 65.7 | 53.1 | 80.9 | 62.7 | 63.0 | 48.8 | 50.8 | 69.2 | 71.2 | ||||||
+ Ours w/ Geo. Prior | 31.1 | 57.9 | 28.3 | 14.5 | 38.5 | 52.7 | 66.3 | 53.6 | 80.7 | 62.5 | 62.8 | 48.1 | 52.9 | 68.4 | 71.4 | ||||||
+ Ours w/ Inst. | 30.7 | 57.2 | 27.9 | 14.5 | 38.4 | 52.8 | 66.4 | 53.3 | 80.8 | 62.4 | 63.7 | 47.7 | 51.7 | 68.6 | 71.1 |
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl | person | rider | car | truck | bus |
bike |
FRCNN | 25.4 | 49.6 | 22.5 | 12.2 | 30.3 | 44.2 | 59.3 | 38.8 | 76.5 | 53.2 | 54.7 | 43.1 | 45.6 | 56.4 | 68.7 | ||||||
+ Ours w/ Sta. Prior | 26.0 | 50.2 | 22.9 | 11.9 | 31.2 | 44.5 | 59.2 | 38.8 | 76.7 | 54.5 | 55.7 | 45.5 | 46.0 | 56.6 | 69.2 | ||||||
+ Ours w/ Geo. Prior | 25.9 | 50.3 | 22.8 | 12.0 | 31.0 | 44.8 | 59.4 | 41.1 | 76.7 | 53.7 | 56.1 | 42.7 | 46.9 | 56.7 | 69.3 | ||||||
+ Ours w/ Inst. | 25.9 | 50.1 | 22.8 | 12.2 | 31.0 | 43.7 | 59.7 | 37.7 | 76.6 | 54.4 | 56.9 | 43.7 | 46.0 | 56.5 | 69.1 |
Appendix D Additional Domain Adaptations
Cityscapes Foggy Cityscapes Table 18 shows the result for domain adaptative object detection from Cityscapes Foggy Cityscapes.
Method | mIoU | road |
building | wall | fence | pole |
terrain | sky | person | rider | car | truck | bus | train |
bike | ||||||||||
DAFormer | 74.7 | 97.8 | 83.0 | 88.7 | 60.4 | 59.4 | 57.4 | 60.7 | 72.8 | 89.2 | 64.9 | 81.2 | 76.6 | 54.4 | 93.2 | 75.9 | 86.9 | 80.7 | 62.8 | 73.8 | ||||||||||
+ Ours w/ Inst. | 75.5 | 98.0 | 84.2 | 89.7 | 59.7 | 61.9 | 57.7 | 60.4 | 73.3 | 89.4 | 62.2 | 82.4 | 76.7 | 56.3 | 92.8 | 82.2 | 88.0 | 82.0 | 63.5 | 73.9 |
Appendix E Additional Analysis
Pixel Accuracy Visualizations Figure 11 show per-pixel accuracy visualization. For both Cityscapes [4] and ACDC [32], a noticeable predominance of red over blue is observed, showing a clear advantage of our method upon DAFormer [12] for semantic segmentation.
BDD100K Clear Rainy Object Detection. Qualitative comparisons are shown in Figure 12. Our method demonstrates superior object detection under rainy conditions by accurately identifying vehicles and minimizing false positives, such as the misidentified pedestrians and cars evident in the 2PCNet [17] predictions.
Cityscapes DarkZurich Semantic Segmentation. Qualitative comparisons are shown in Figure 13. The proposed method exhibits a segmentation output that is more closely aligned with groundtruth, particularly while predicting road boundaries and consistently identification of urban elements such as sidewalk, terrain, and traffic signs. This indicates that our method has superior domain adaptation capabilities in challenging low-light conditions.
Appendix F Additional Technical Details
Ground Truth Segmentation to Boxes: Instance-based image warping requires bounding boxes, which are not provided in some semantic segmentation benchmarks like GTA [28]. To circumvent this, we generate “from-seg” bounding boxes from ground truth semantic segmentation maps. We do this by identifying connected components that represent individual instances of foreground categories — specifically, traffic lights, traffic signs, persons, riders, cars, trucks, buses, trains, motorcycles, and bikes.
For each “from-seg” instance, we compute the bounding boxes by finding the minimum enclosing axis-aligned rectangle. These “from-seg” bounding boxes are then employed for instance-level image warping in a similar manner to ground truth bounding boxes.